OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. Tawny found $1.95 in quarters and dimes in the cushions of her couch. There were four times as many dimes as there were quarters. How many dimes and quarters did she find?

OpenStudy (anonymous):

q = # of quarters d = # of dimes .25q + .1d = 1.95 d = 4q since there are 4 times as many dimes as there are quarters. use substitution: .25q + .1(4q) = 1.95 .25q + .4q = 1.95 .65q = 1.95 q = 3 so d = 12

OpenStudy (thewayiam983):

L: Let x = # of quarters worth 25 cents each Let 4x = # of dimes worth 10 cents each E: 195 cents = 25(x) + 10(4x) S: 195 = 65x 3 = x C: 195 = 25(3) + 10(12) 195 = 75 + 120 195 = 195 A: There are 3 quarters and 12 dimes

OpenStudy (anonymous):

great thanx sis and 983

OpenStudy (anonymous):

welcome